OpenStudy (anonymous):

simplify: x^2 - 16x + 64 over x^2 - 64

OpenStudy (wasiqss):

lolx easiest

OpenStudy (wasiqss):

answer is (x-8) over (x+8)

OpenStudy (lalaly):

\[\frac{x^2-16x+64}{x^2-64}= \frac{ \cancel{(x-8)}(x-8)}{\cancel{(x-8)}(x+8)}\]

OpenStudy (wasiqss):

wow lalaly ur work is soo organised :P

OpenStudy (anonymous):

Thank you all! :D Sorry, I'm not too good at algebra...

OpenStudy (lalaly):

=)

OpenStudy (wasiqss):

honey u r gud at nthin :D

OpenStudy (anonymous):

haha i know no need for the reminder

OpenStudy (wasiqss):

hehe

OpenStudy (anonymous):

lol maybe i shouldn't have given you a medal hon :P

OpenStudy (wasiqss):

haha u have given now so no choice back :D

OpenStudy (anonymous):

i know >.<